OpenStudy (evan7):

i need help with multiplication and division equations.. ex: 68 = r/4

OpenStudy (maddy1251):

@evan7 not to fear. For the equation \[68= \frac{ r }{ 4 }\]

OpenStudy (maddy1251):

We want to isolate the variable to solve for it, since that is the desirable outcome we are looking for. For this particular instant, do the opposite to the variable as the operation that is being performed upon it. In this case, multiply '4' to both sides.

OpenStudy (maddy1251):

\[68 x 4 = \frac{ r }{ 4 } x 4\]

OpenStudy (maddy1251):

\[272=r\]

OpenStudy (maddy1251):

Now, if this was in the form of \[64=4r\]

OpenStudy (maddy1251):

We would divide each side by '4' since division is the opposite of multiplication, that's how we will get 'r' alone on one side.

OpenStudy (maddy1251):

\[64/4=4r/4\]

OpenStudy (maddy1251):

\[r=16\]

OpenStudy (evan7):

thank you!!!!!.. just this helped me out.. i did all my homework .. well except for one question cus it confused me too

OpenStudy (evan7):

a cat can jump the height of up to 5 times the length of its tail. how high can a cat jump if its tail is 13 inches long?

OpenStudy (evan7):

is it 5x=13 .. x =2.5?

OpenStudy (maddy1251):

My calculator says x=2.6, but you where just a tenth off. :-)